# TabMGP: Martingale Posterior with TabPFN

### v1

@@ 4 Related works
Recent analyses (Falck et al., 2024; Ye and Namkoong, 2024; Nagler and Rügamer, 2025) show that the predictive rules induced by foundation models, such as TabPFN, are not martingales. Building on this line of work, we take the further step of testing the weaker a.c.i.d. condition (Battiston and Cappello, 2025) and show that TabPFN fails to satisfy it as well (Appendix C). Nevertheless, we find that the resulting TabMGP exhibits convergence (Appendix F) and near-nominal coverage (Section 5) in practice, suggesting that the usual sufficient conditions for validity are stricter than necessary.
Interestingly, Nagler and Rügamer (2025) adopt an approach closely aligned with ours, employing TabPFN as the predictive rule in the martingale posterior framework. Their finding that TabPFN is not a martingale leads them to abandon it as a predictive rule after the initial step, instead using its output to initialise a copula-based update that preserves martingale validity. By contrast, we retain TabPFN as the predictive rule throughout and interpret its failure to satisfy both the martingale and a.c.i.d. conditions as evidence that validity can be achieved under weaker assumptions.
@@
Bayesian boostrap
The Bayesian bootstrap (Rubin, 1981) is a predictive rule where future observations $Z_{i+1}$ are uniformly drawn from the observations up to $i$:
@@
with $F_{i}$ the empirical distribution of $z_{1:i}$. As this rule is discrete and unsmoothed, its use is recommended only when the functional of interest arises from a parametric likelihood loss, for which the smoothness of $F_{\infty}$ is irrelevant (Fong et al., 2023; Dellaporta et al., 2022).
Newton’s algorithm.
Rooted in Bayesian nonparametrics, Newton’s algorithm (Newton et al., 1998) was initially developed for unsupervised sequential learning in mixture models. It was subsequently re-interpreted as a predictive rule that generates a c.i.d. sequence of observations (Fortini and Petrone, 2020). This approach, however, is not suitable for high-dimensional $\theta$, as it requires numerical integration of an intractable normalising constant.
Copula-based updates.
For applications requiring smoothness of $F_{\infty}$, the mainstream implementation of the martingale posterior is dominated by copula-based updates. Introduced by Hahn et al. (2018) and adopted as the default predictive rule in Fong et al. (2023), these updates smooth $F_{N}$ by attaching a bivariate copula to each new observation while maintaining the martingale property on the predictive rule exactly. Subsequent work has extended this framework with autoregressive Gaussian-process likelihoods for higher-dimensional features (Ghalebikesabi et al., 2023), direct quantile updates (Fong and Yiu, 2024b), log-concave shape constraints (Cui and Walker, 2024), and vine copulas for very high dimensional features (Huk et al., 2024). All these variants preserve the exact martingale guarantee of the copula framework, but each introduces at least one smoothing or dependence hyperparameter, whose value controls the spread of the resulting martingale posterior.
Parametric plug-in predictives.
For parametric inference, Walker (2022) and Holmes and Walker (2023) proposed a predictive rule of the form $Z_{i+1}\sim p(\cdot\mid\widehat{\theta}_{i})$, where $p(\cdot\mid\theta)$ is an assumed parametric model and $\widehat{\theta}_{i}$ is a point estimate (e.g., maximum likelihood estimate) computed from all previous observations $z_{1:i}$. The asymptotic properties of this predictive rule are studied in Fong and Yiu (2024a). However, it requires that the assumed parametric model provides a reasonable approximation to the data distribution in order to yield sensible uncertainty quantification.
Permutation-equivariant neural network.
Martingale posteriors have also been explored in the context of neural processes (Lee et al., 2023), where a permutation-equivariant neural network is used to construct a martingale posterior over latent variables. While this approach demonstrates the potential of using neural networks within a martingale posterior framework, the focus is still on prediction instead of inference, and it remains unclear how to generalize to broader statistical inference tasks.
@@ 5 Experiments
We compare TabMGP with other martingale posterior constructions using synthetic and real-world data. Our experiments focus on small- to moderate-$n$ settings where the prior specification of the classical Bayesian posterior is important, making the martingale posterior particularly well suited.
@@
$\Pr_{z_{1:n}\sim F_{0}}\bigl(\theta(F_{0})\in C_{1-\alpha}(z_{1:n})\bigr),$ (4)
namely, the proportion which the credible set contains the population risk minimiser, $\theta(F_{0})=\arg\min_{\theta}\int\ell(z,\theta)\,\mathrm{d}F_{0}(z)$, under repeated draws of $z_{1:n}$ from the true data-generating distribution $F_{0}$. Although frequentist in nature, this metric is commonly used for comparing Bayesian posteriors. We view higher coverage (closer to the nominal level), paired with a reasonably sized credible set, as indicative of more reliable and meaningful uncertainty quantification for the functional of interest.
Posterior constructions.
We compare TabMGP against the martingale posteriors suggested in Fong et al. (2023): Bayesian bootstrap (BB) and the bivariate copula update (Copula). As a baseline, we also include the classical Bayesian posterior (Bayes) with an appropriate model and an isotropic Gaussian prior with variance $10^{2}$; the model choice is discussed later in this section when introducing the loss functions for this experiment. The exact computation details of the posteriors are provided in Appendix E.1.
Loss functions.
We use negative log-likelihood as the loss in (1), and the corresponding model in the Bayes posterior. For continuous response (linear regression), we use $\ell(x,y,\theta)=(y-[1\ x^{\top}]\theta)^{2}$ which is the negative log-likelihood of ${\mathcal{N}}([1\ x^{\top}]\theta,1)$. For $K$-class categorical response (logistic regression), we use $\ell(x,y,\theta)=-\log\Pr(y=k)$, where $\Pr(y=k)$ is the $k$-th entry of the softmax applied to the logits $([1\ x^{\top}]\theta_{1},\ldots,[1\ x^{\top}]\theta_{K})$, and $\theta_{1},\ldots,\theta_{K}$ are the class-specific coefficients. We also set appropriate constraints on the coefficients to ensure identifiability in the loss function; see Appendix E.2 for details.
Data-generating setups.
We describe the true data-generating distribution $F_{0}$ of this experiment. For each $F_{0}$, we repeatedly sample 100 repetitions $\{z_{1:n}^{(r)}\}_{r=1}^{100}$ from $F_{0}$. We refer to $F_{0}$ together with the procedure used to generate $z_{1:n}$ as a setup. The $n$ is chosen to be as small as possible without rendering the design matrix ill-conditioned. In total, we have 30 setups (11 synthetic, 19 real-data) in this experiment.
Let $\beta_{0}$ be a set of regression coefficients and $x_{i}$ i.i.d. samples from a hypercube $1,1]^{10}$. For the synthetic setups, we generate $y\mid x$ with:
@@
    $y_{i}=x^{\top}_{i}\beta_{0}+\epsilon_{i}$ where $\epsilon_{i}\overset{iid}{\sim}{\mathcal{N}}(0,1)$;
@@
    $y_{i}=x^{\top}_{i}\beta_{0}+\epsilon_{i}$ where $\epsilon_{i}$ are i.i.d. from $t_{5}$, $t_{4}$ or $t_{3}$;
@@
    $y_{i}=x^{\top}_{i}\beta_{0}+\epsilon_{i}$ where $\epsilon_{i}$ is heteroscedastic and depends on $x_{i}$. We have three different strength of heteroscedasticity: $s_{1}$, $s_{2}$, $s_{3}$, from weak to strong;
@@
    $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta_{0}))$, with a logistic link function $L$;
@@
    $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta_{0}))$, where $L$ is the distribution function of a Gaussian mixture model (GMM). We have models with three different location parameters: GMM(0)(0), GMM(1)(-1), GMM(2)(-2).
For the 19 real-data setups, we use datasets from the OpenML and UCI repositories (9 continuous response, 8 binary response, and 2 multiclass categorical response). As $F_{0}$ is unknown in the real-data setups, we take the empirical distribution of the entire dataset as the truth $\widetilde{F}_{0}$. The details of all data-generating setups are given in Appendix E.3.
Metrics for comparing posteriors.
For each posterior, we construct joint credible sets $\widehat{C}_{1-\alpha}(z_{1:n})$ and compare their coverage (4) and ‘size’. Our approach of constructing $\widehat{C}_{1-\alpha}(z_{1:n})$ is described in Appendix E.4. We consider a posterior more reliable and meaningful when its credible set is small while achieving coverage at least at the the nominal level $1-\alpha$.
In the synthetic setups, the coverage is given by $\frac{1}{100}\sum^{100}_{r=1}{\bm{1}}\bigl[\theta(F_{0})\in\widehat{C}_{1-\alpha}(z^{(r)}_{1:n})\bigr$], where $\{z_{1:n}^{(r)}\}_{r=1}^{100}$ are 100 replicates of $z_{1:n}$ drawn from $F_{0}$. In the real-data setups, as $F_{0}$ is unknown, we take the empirical distribution of the entire dataset $\mathcal{D}$} as the truth $\widetilde{F}_{0}$, and compute the coverage on $\theta(\widetilde{F}_{0})=\arg\min_{\theta}\sum_{z\in{\mathcal{D}}}\ell(z,\theta)$. That is, the coverage is $\frac{1}{100}\sum^{100}_{r=1}{\bm{1}}\bigl[\theta(\widetilde{F}_{0})\in\widehat{C}_{1-\alpha}(z^{(r)}_{1:n})\bigr$], where $z^{(r)}_{1:n}$ denotes the $r$-th random draws (without replacement) from $\mathcal{D}$}; see Appendix E.3 for details. To quantify the ‘size’ of a credible set, we use the trace of the posterior covariance matrix as a proxy.
@@ 5.1 Results
Convergence of $\theta(F_{N})$.
To conduct statistical inference, we require $\theta(F_{N})$ to converge $\mathbb{P}$}-a.s. as $N\to\infty$. This convergence is commonly assessed via trace plots that track each dimension of $\theta(F_{N})$ as $N$ increases, but this quickly becomes impractical for high-dimensional $\theta$. Instead, we track the expected $L_{1}$-norm of $\theta(F_{N})$ relative to the empirical risk minimiser $\theta(F_{n})$ as $N$ increases; see Appendix F for the exact definition. We present the trajectories in Figure 1, where each trajectory corresponds to one of the 30 setups. This figure suggests that the trajectories (and hence $\theta(F_{N})$) generally converge as $N$ increases, albeit with more noise in some setups.
@@
We present the marginals of the posteriors for a realisation of $z_{1:n}$ in Figure 2 (intercept of ‘concrete’), with the remaining dimension and setups provided in Appendix G. In general, the posterior means are similar across methods, but TabMGP often exhibits skewness and multimodal structure compared with the other posteriors, which are typically Gaussian-like. This suggests that TabMGP is more expressive than the other posteriors.
@@
Figure 2: Posterior densities of the intercept in the ‘concrete’ setup. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean as dots. The vertical dotted red and blue lines correspond to $\theta(F_{0})$ and $\theta(F_{n})$, respectively.
Joint credible sets of linear regression.
@@
The results for synthetic data (top and middle) and real-data (bottom) setups are shown in Table 1. In the real-data setups, the rows are ordered by the $n/p$ ratio in descending order, with $p\coloneq\dim(\theta)$.
@@
TabMGP is competitive in both coverage and size (values in parentheses) of the credible set, performing particularly well on real data. Notably, it can attain almost 100% coverage in many setups while maintaining a small credible set. Similarly, Copula provides good coverage and small credible sets, although its sets are generally larger than those of TabMGP in most real-data setups. Bayes achieves close to 100% coverage in all setups, but its credible sets are much larger than those of TabMGP due to its diffuse prior. In contrast, BB tends to produce overly small sets with poor coverage in the synthetic setups (top and middle) and in low $n/p$ settings (last few rows of Table 1). This reflects limited variability in forward samples when bootstrapping from a small $z_{1:n}$ (e.g., $n=20$ in the synthetic data setups).
$z_{1:n}$ TabMGP $B$ Copula Bayes
${\mathcal{N}}(0,1)$ 1.00 (0.45) 0.55 (0.09) 0.99 (0.35) 1.00 (1.29)
$t_{5}$ 1.00 (0.49) 0.65 (0.11) 0.99 (0.37) 1.00 (1.29)
$t_{4}$ 0.99 (0.51) 0.64 (0.12) 0.98 (0.37) 1.00 (1.30)
$t_{3}$ 1.00 (0.48) 0.66 (0.14) 0.97 (0.35) 0.98 (1.30)
$s_{1}$ 1.00 (0.37) 0.50 (0.03) 1.00 (0.36) 1.00 (1.30)
$s_{2}$ 1.00 (0.36) 0.46 (0.02) 1.00 (0.36) 1.00 (1.30)
$s_{3}$ 1.00 (0.33) 0.53 (0.02) 1.00 (0.37) 1.00 (1.30)
concrete 0.91 (0.06) 0.80 (0.05) 1.00 (0.12) 1.00 (0.10)
quake 1.00 (0.03) 0.95 (0.07) 1.00 (0.18) 0.98 (0.09)
airfoil 0.96 (0.08) 0.93 (0.05) 0.97 (0.11) 1.00 (0.12)
energy 1.00 (0.04) 0.80 (0.01) 1.00 (0.06) 1.00 (0.14)
fish 0.90 (0.10) 0.79 (0.07) 0.95 (0.10) 1.00 (0.16)
kin8nm 0.95 (0.13) 0.77 (0.09) 0.63 (0.08) 0.99 (0.22)
auction 0.99 (0.66) 0.86 (0.35) 0.97 (1.44) 1.00 (1.64)
grid 0.97 (0.13) 0.72 (0.08) 0.88 (0.10) 1.00 (0.31)
abalone 0.97 (0.95) 0.66 (0.69) 0.94 (0.98) 0.98 (1.62)
@@
Table 1: Experiments on linear regression: i.i.d. errors (top), dependent errors (middle), and real data (bottom). We compare TabMGP (ours), BB, Copula, and Bayes. Reported values are coverage and the trace of the posterior covariance (median, in parentheses), computed over 100 repetitions. Target coverage is 0.95.
Joint credible sets of logistic regression.
@@
The results for synthetic (top) and real-data (middle and bottom) setups are reported in Table 2. Copula was not applied to multinomial logistic regression (bottom) as the implementation of Fong et al. (2023) is not directly applicable. In the real-data setups, the rows are ordered by the $n/p$ ratio in descending order within each section.
@@
TabMGP generally yields the smallest credible sets with good coverage (typically above 95%), except in the synthetic setups, where it slightly under-covers. Bayes is the next best, attaining nearly 100% coverage and producing credible sets that are generally smaller than those of BB and Copula, but larger than those of TabMGP. BB performs reasonably well in binary logistic regression (top and middle), although its credible sets are larger than those of TabMGP. However, BB severely under-covers in multinomial logistic regression (bottom), where $p$ is much greater than in the other logistic regression setups. Copula performs poorly in logistic regression, producing trivial credible sets. This indicates that Copula is sensitive to the choice of bandwidth, as the same value yields good performance in the linear regression experiment.
$z_{1:n}$ TabMGP $B$ Copula Bayes
Logistic 0.84 (1.30) 0.83 (1.96) 1.00 (2.72) 0.99 (1.94)
GMM(0)(0) 0.91 (2.44) 0.91 (4.14) 1.00 (4.01) 0.99 (3.66)
GMM(1)(-1) 0.83 (1.27) 0.86 (1.95) 1.00 (2.64) 1.00 (1.95)
GMM(2)(-2) 0.78 (0.88) 0.78 (1.36) 1.00 (2.08) 1.00 (1.37)
rice 1.00 (3.74) 0.88 (7.35) 1.00 (28.4) 1.00 (6.66)
sepsis 0.95 (0.69) 1.00 (2.42) 1.00 (10.0) 1.00 (3.23)
banknote 0.99 (1.16) 1.00 (1.48) 1.00 (1.53) 1.00 (1.48)
mozilla 0.96 (1.42) 0.97 (2.04) 1.00 (2.68) 0.96 (1.42)
skin 1.00 (1.39) 1.00 (1.84) 1.00 (1.80) 1.00 (1.49)
blood 0.87 (1.07) 0.95 (2.51) 1.00 (5.40) 1.00 (2.06)
phoneme 1.00 (1.32) 0.95 (1.92) 1.00 (5.62) 1.00 (2.15)
telescope 0.99 (5.96) 0.99 (10.7) 1.00 (50.7) 1.00 (9.39)
yeast 0.97 (11.1) 0.68 (25.1) 1.00 (26.9)
wine 0.93 (11.0) 0.79 (39.7) 0.98 (33.1)

### v2

@@ 4 Related works
Recent analyses (Falck et al., 2024; Ye and Namkoong, 2024; Nagler and Rügamer, 2025) show that the predictive rules induced by foundation models, such as TabPFN, operate outside the traditional martingale property. Building on this line of work, we take the further step of assessing the more flexible almost conditionally identically distributed (a.c.i.d.) condition (Battiston and Cappello, 2025) and observe that TabPFN departs from this requirement as well. Nevertheless, we find that the resulting TabMGP exhibits consistent path stability and near-nominal coverage in practice. This suggests that the usual sufficient conditions are stricter than necessary and that validity can be achieved under a more relaxed set of assumptions than previously theorised.
Interestingly, Nagler and Rügamer (2025) adopt an approach closely aligned with ours, employing TabPFN as the predictive rule in the martingale posterior framework. Their finding that TabPFN is not a martingale leads them to abandon it as a predictive rule after the initial step, instead using its output to initialise a copula-based update that preserves martingale validity. By contrast, we retain TabPFN as the predictive rule throughout and interpret its departure from both the martingale and a.c.i.d. conditions as evidence that validity can be achieved under weaker assumptions.
A concurrent effort in Fortini et al. (2026) focuses on posterior inference for $F_{\infty}\mid z_{1:n}$ using a predictive central limit theorem (CLT). However, their framework does not incorporate the generalised Bayes flavour of inference for a parameter $\theta$ defined by a user-specified loss function. The predictive CLT has distinct advantages in supervised settings, and extending the predictive CLT to functionals of $\theta(F_{\infty})$ would make for interesting future work.
@@
Bayesian boostrap The Bayesian bootstrap (Rubin, 1981) is a predictive rule where future observations $Z_{i+1}$ are uniformly drawn from the observations up to $i$:
@@
with $F_{i}$ the empirical distribution of $z_{1:i}$. As this rule is discrete, its use is recommended only when the functional of interest arises from a parametric likelihood loss, for which the smoothness of $F_{\infty}$ is irrelevant (Fong et al., 2023; Dellaporta et al., 2022).
Newton’s algorithm. Rooted in Bayesian nonparametrics, Newton’s algorithm (Newton et al., 1998) was initially developed for unsupervised sequential learning in mixture models. It was subsequently re-interpreted as a predictive rule that generates a c.i.d. sequence of observations (Fortini and Petrone, 2020). This approach, however, is not suitable for high-dimensional $\theta$, as it requires numerical integration of an intractable normalising constant.
Copula-based updates. For applications requiring smoothness of $F_{\infty}$, the mainstream implementation of the martingale posterior is dominated by copula-based updates. Introduced by Hahn et al. (2018) and adopted as the default predictive rule in Fong et al. (2023), these updates smooth $F_{N}$ by attaching a bivariate copula to each new observation while maintaining the martingale property on the predictive rule exactly. Subsequent work has extended this framework with autoregressive Gaussian-process likelihoods for higher-dimensional features (Ghalebikesabi et al., 2023), direct quantile updates (Fong and Yiu, 2024b), log-concave shape constraints (Cui and Walker, 2024), and vine copulas for very high dimensional features (Huk et al., 2024). All these variants preserve the exact martingale guarantee of the copula framework, but each introduces at least one smoothing or dependence hyperparameter, whose value controls the spread of the resulting posterior. These hyperparameters are generally difficult to tune and must be retuned for each new dataset.
Parametric plug-in predictives. For parametric inference, Walker (2022) and Holmes and Walker (2023) proposed a predictive rule of the form $Z_{i+1}\sim p(\cdot\mid\widehat{\theta}_{i})$, where $p(\cdot\mid\theta)$ is an assumed parametric model and $\widehat{\theta}_{i}$ is a point estimate (e.g., maximum likelihood estimate) computed from all previous observations $z_{1:i}$. The asymptotic properties of this predictive rule are studied in Fong and Yiu (2024a). However, it requires that the assumed parametric model provides a reasonable approximation to the data distribution in order to yield sensible uncertainty quantification.
Permutation-equivariant neural network. Martingale posteriors have also been explored in the context of neural processes (Lee et al., 2023), where a permutation-equivariant neural network is used to construct a martingale posterior over latent variables. While this approach demonstrates the potential of using neural networks within a martingale posterior framework, the focus is still on prediction instead of inference, and it remains unclear how to generalise to broader statistical inference tasks.
@@ 5 Experiments
We compare TabMGP with other martingale posterior constructions using synthetic and real-world data. Our experiments focus on small- to moderate-$n$ settings where the prior specification of the classical Bayesian posterior is important, making the MGP particularly well suited.
@@
$\Pr_{z_{1:n}\sim F^{\star}}\bigl(\theta(F^{\star})\in C_{1-\alpha}(z_{1:n})\bigr),$ (4)
namely, the proportion which the credible set contains the population risk minimiser, $\theta(F^{\star})=\arg\min_{\theta}\int\ell(z,\theta)\,\mathrm{d}F^{\star}(z)$, under repeated draws of $z_{1:n}$ from the true data-generating distribution $F^{\star}$. Although frequentist in nature, this metric is commonly used for comparing Bayesian posteriors.
Posterior constructions. We compare TabMGP against the martingale posteriors suggested in Fong et al. (2023): Bayesian bootstrap (BB) and the bivariate copula update (Copula). As a baseline, we also include the classical Bayesian posterior (Bayes) using asymptotic approximation of the posterior as the prior. We also compare TabMGP against this asymptotic approximation of the posterior (Asymptotic). The model choice is discussed later in this section when introducing the loss functions for this experiment. The exact computation details of the posteriors are provided in Appendix E.1.
Loss functions. We use negative log-likelihood as the loss in (1), and the corresponding model in the Bayes posterior. For continuous response (linear regression), we use $\ell(x,y,\theta)=(y-[1\ x^{\top}]\theta)^{2}$ which is the negative log-likelihood of ${\mathcal{N}}([1\ x^{\top}]\theta,1)$. For $K$-class categorical response (logistic regression), we use $\ell(x,y,\theta)=-\log\Pr(y=k)$, where $\Pr(y=k)$ is the $k$-th entry of the softmax applied to the logits $([1\ x^{\top}]\theta_{1},\ldots,[1\ x^{\top}]\theta_{K})$, and $\theta_{1},\ldots,\theta_{K}$ are the class-specific coefficients. We also set appropriate constraints on the coefficients to ensure identifiability in the loss function; see Appendix E.2 for details.
Data-generating setups. We describe the true data-generating distribution $F^{\star}$ of this experiment. For each $F^{\star}$, we repeatedly sample 100 repetitions $\{z_{1:n}^{(r)}\}_{r=1}^{100}$ from $F^{\star}$. We refer to $F^{\star}$ together with the procedure used to generate $z_{1:n}$ as a setup. The $n$ is chosen to be as small as possible without rendering the design matrix ill-conditioned. In total, we have 30 setups (11 synthetic, 19 real-data).
Let $\beta^{\star}$ be a set of regression coefficients and $x_{i}$ i.i.d. samples from a hypercube $1,1]^{10}$. For the synthetic setups, we generate $y\mid x$ with:
@@
    $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i}$ where $\epsilon_{i}\overset{iid}{\sim}{\mathcal{N}}(0,1)$;
@@
    $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i}$ where $\epsilon_{i}$ are i.i.d. and follows a Student’s t-distribution: $t_{5}$, $t_{4}$ or $t_{3}$;
@@
    $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i}$ where $\epsilon_{i}$ is heteroscedastic and depends on $x_{i}$. We have three different strength of heteroscedasticity: $s_{1}$, $s_{2}$, $s_{3}$, from weak to strong;
@@
    $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta^{\star}))$, with a logistic link function $L$;
@@
    $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta^{\star}))$, where $L$ is the distribution function of a Gaussian mixture model (GMM). We consider three choices of location parameters: GMM(0)(0), GMM(1)(-1), and GMM(2)(-2), where the argument denotes the mean of one mixture component.
For the 19 real-data setups, we use datasets from the OpenML and UCI repositories (9 continuous response, 8 binary response, and 2 multiclass categorical response). As $F^{\star}$ is unknown in the real-data setups, we take the empirical distribution of the entire dataset as the truth $\widetilde{F}^{\star}$. The details of all data-generating setups are given in Appendix E.3.
Metrics for comparing posteriors. For each posterior, we construct joint credible sets $\widehat{C}_{1-\alpha}(z_{1:n})$ and compare their coverage (4) and ‘size’. Our approach of constructing $\widehat{C}_{1-\alpha}(z_{1:n})$ is described in Appendix E.4. We consider a posterior more reliable and meaningful when its credible set is small while achieving coverage at least $1-\alpha$.
In the synthetic setups, the coverage is given by $\frac{1}{100}\sum^{100}_{r=1}{\bm{1}}\bigl[\theta(F^{\star})\in\widehat{C}_{1-\alpha}(z^{(r)}_{1:n})\bigr$], where $\{z_{1:n}^{(r)}\}_{r=1}^{100}$ are 100 replicates of $z_{1:n}$ drawn from $F^{\star}$. In the real-data setups, as $F^{\star}$ is unknown, we take the empirical distribution of the entire dataset $\mathcal{D}$} as the truth $\widetilde{F}^{\star}$, and compute the coverage on $\theta(\widetilde{F}^{\star})=\arg\min_{\theta}\sum_{z\in{\mathcal{D}}}\ell(z,\theta)$. That is, the coverage is $\frac{1}{100}\sum^{100}_{r=1}{\bm{1}}\bigl[\theta(\widetilde{F}^{\star})\in\widehat{C}_{1-\alpha}(z^{(r)}_{1:n})\bigr$], where $z^{(r)}_{1:n}$ denotes the $r$-th random draws (without replacement) from $\mathcal{D}$}; see Appendix E.3 for details. To quantify the ‘size’ of a credible set, we use the trace of the posterior covariance matrix as a proxy.
@@ 5.1 Results
Convergence of $\theta(F_{N})$. To conduct statistical inference, we require $\theta(F_{N})$ to converge $\mathbb{P}$}-a.s. as $N\to\infty$. This convergence is commonly assessed using trace plots that track each component of $\theta(F_{N})$ as $N$ increases, but this approach quickly becomes impractical in high dimensions. Instead, we track the expected $L_{1}$-norm of $\theta(F_{N})$ relative to the empirical risk minimiser $\theta(F_{n})$ as $N$ increases; see Appendix F for the precise definition. For convergent $\theta(F_{N})$, we desire to see the expected $L_{1}$-norm stabilise at a non-zero constant as $N$ increases.
We present the trajectories in Figure 3, where each trajectory corresponds to one of the 30 setups. For completeness, we also report the $L_{1}$-norm between each realisation $F_{N}$ and $F_{n}$ in Figure 6. In general, we observe that $N=500$ is sufficient for convergence in most setups, while some setups are near convergence and are expected to converge as the number of forward sampling steps increases.
@@
Shape of the posteriors. We present the marginals of the posteriors for a realisation of $z_{1:n}$ in Figure 4 (intercept of the real-data ‘concrete’ setup), with the remaining dimension and setups provided in Appendix G. In general, the posterior means are similar across methods, but TabMGP often exhibits skewness and multimodal structure compared with the other posteriors, which are typically Gaussian-like. This suggests that TabMGP is more expressive.
@@
Figure 4: Posterior densities of the intercept in the ‘concrete’ setup. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
@@
Joint credible sets of linear regression. The results for synthetic data (top and middle) and real-data (bottom) setups are shown in Table 1. In the real-data, the rows are ordered by the $n/p$ ratio in descending order, with $p\coloneq\dim(\theta)$.
@@
TabMGP is competitive in both coverage and credible set size, performing particularly well on real data. Notably, it can attain nearly 100% coverage in many setups while maintaining small credible sets. Similarly, Copula provides good coverage with small credible sets, although its performance is less consistent across real-data setups (e.g., severe undercoverage in kin8nm and overcoverage in quake). This is expected, as bespoke Gaussian-copula-based predictive rules have a structural advantage over TabMGP in Gaussian (or near-Gaussian) setups. Bayes and Asymptotic produce excessively large credible sets in all synthetic setups, as both rely on asymptotic approximations that are less accurate in low-$n$ settings. They are also more susceptible to model misspecification and tend to underperform relative to TabMGP. In contrast, BB tends to produce overly small sets with poor coverage in the synthetic setups (top and middle) and in low $n/p$ settings (last few rows of Table 1). This reflects limited variability in forward samples when bootstrapping from a small $z_{1:n}$ (e.g., $n=20$ in the synthetic data setups).
@@
Table 1: Experiments on linear regression: i.i.d. errors (top), dependent errors (middle), and real data (bottom). We compare TabMGP (ours), BB, Copula, Bayes and Asymptotic. The rows corresponding to real data are sorted by $n/p$ ratio, i.e, bottom rows are harder. Reported values are coverage (Rate) and the median of trace of the posterior covariance (Size), computed over 100 repetitions. Target coverage is 0.95.
TabMGP $B$ Copula Bayes Asymptotic
Setup Rate Size Rate Size Rate Size Rate Size Rate Size
${\mathcal{N}}(0,1)$ 1.00 0.45 0.55 0.09 0.99 0.35 1.00 0.65 1.00 1.31
$t_{5}$ 1.00 0.49 0.65 0.11 0.99 0.37 1.00 0.65 1.00 1.31
$t_{4}$ 0.99 0.51 0.64 0.12 0.98 0.37 0.99 0.65 1.00 1.31
$t_{3}$ 1.00 0.48 0.66 0.14 0.97 0.35 0.98 0.65 0.98 1.31
$s_{1}$ 1.00 0.37 0.50 0.03 1.00 0.36 1.00 0.65 1.00 1.31
$s_{2}$ 1.00 0.36 0.46 0.02 1.00 0.36 1.00 0.65 1.00 1.31
$s_{3}$ 1.00 0.33 0.53 0.02 1.00 0.37 1.00 0.65 1.00 1.31
concrete 0.91 0.06 0.80 0.05 1.00 0.12 0.87 0.05 1.00 0.10
quake 1.00 0.03 0.95 0.07 1.00 0.18 0.91 0.04 0.98 0.09
airfoil 0.96 0.08 0.93 0.05 0.97 0.11 0.96 0.06 1.00 0.12
energy 1.00 0.04 0.80 0.01 1.00 0.06 1.00 0.07 1.00 0.14
fish 0.90 0.10 0.79 0.07 0.95 0.10 0.91 0.08 1.00 0.16
kin8nm 0.95 0.13 0.77 0.09 0.63 0.08 0.84 0.11 0.99 0.22
auction 0.99 0.66 0.86 0.35 0.97 1.44 1.00 0.81 1.00 1.65
grid 0.97 0.13 0.72 0.08 0.88 0.10 0.97 0.16 1.00 0.31
abalone 0.97 0.95 0.66 0.69 0.94 0.98 0.84 0.81 0.98 1.64
@@
Joint credible sets of logistic regression. The results for synthetic (top) and real-data (middle and bottom) setups are in Table 2. Copula was not applied to multinomial logistic regression (bottom) as the implementation of Fong et al. (2023) is not directly applicable. In the real-data setups, the rows are ordered by decreasing $n/p$ within each section.
@@
In the synthetic setups, the Asymptotic credible sets achieve the best coverage and size, as expected, since the asymptotic approximation is reasonably accurate at the larger sample size ($n=100$). TabMGP, however, shows the most consistent performance in the real-data setups, often achieving good coverage with the smallest credible sets. TabMGP performs particularly well relative to competing methods in the ‘harder’ setups, i.e., the bottom rows with low $n/p$ ratios. Asymptotic also provides reasonable coverage in the real-data setups, although its credible sets are typically much wider than those of TabMGP. This is likely due to model misspecification, which results in miscalibrated credible sets. BB performs reasonably well in binary logistic regression (top and middle), but its credible sets are larger than those of TabMGP. However, BB severely under-covers in multinomial logistic regression (bottom), where $p$ is much larger than in the other logistic regression setups. Copula performs poorly in logistic regression, producing trivial credible sets. This suggests that Copula is sensitive to the choice of bandwidth, as the same value yields good performance in the linear regression experiment.